OpenStudy (anonymous):

What is the potential difference needed to accelerate a He+ ion (charge +e; 4 amu) from rest to a speed of 2*10^6 m/s?

OpenStudy (anonymous):

Okay, so this should just be an energy problem right? Since it starts from rest\[K=U\] and \[K=\frac{1}{2}mv^{2}\]and \[U=qV\] substitute solve for \[V=\frac{mv^{2}}{2q}\] Correct?

OpenStudy (prakharluv):

Let potential difference = V q V = (1/2)*m*(vf^2 - vi^2) Can you substitute values and calculate V?

OpenStudy (anonymous):

Yes. That is what I did, except when I substituted the values in I had: \[q = +1.602*10^{-19}C\]\[m = 4*1.67*10^{-27} kg= 6.68*10^{-27}kg\]\[v_{f}-v_{i}=0m/s-2*10^{6}m/s=-2*10^6m/s\]Plug those values in I have \[V=\frac{mv^2}{2q}=\frac{(6.68*10^{-27}kg)*(-2*10^6m/s)^{2}}{2*(1.602*10^{-19}C)}=83395.7 V\]...oh my gosh this is embarrassing. I was forgetting to square the delta v term whenever I plugged it in last night. So I kept getting .041 V .. lol. Thanks.